### v3

# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

Abstract
Enhancing low-light images remains a critical challenge in computer vision, as does designing lightweight models for edge devices that can handle the computational demands of deep learning. In this article, we introduce an extended version of the Channel-Prior and Gamma-Estimation Network (CPGA-Net), termed CPGA-Net+, which incorporates an attention mechanism driven by a reformulated Atmospheric Scattering Model and effectively addresses both global and local image processing through Plug-in Attention with gamma correction. These innovations enable CPGA-Net+ to achieve superior performance on image enhancement tasks for supervised and unsupervised learning, surpassing lightweight state-of-the-art methods with high efficiency. Furthermore, we provide a theoretical analysis showing that our approach inherently decomposes the enhancement process into restoration and lightening stages, aligning with the fundamental image degradation model. To further optimize efficiency, we introduce a block simplification technique that reduces computational costs by more than two-thirds. Experimental results validate the effectiveness of CPGA-Net+ and highlight its potential for applications in resource-constrained environments.
Index Terms:
Atmospheric Scattering Model, Low-Light Image Enhancement, Gamma Correction, Channel Prior, Explainable AI.
I Introduction
Low-light image capture, whether indoors or outdoors, poses significant challenges for accurate visual analysis. The limited light reflection often results in degraded image quality, including color inaccuracies and increased noise levels. These issues can significantly affect the performance and reliability of light-sensitive applications, such as transportation surveillance and Advanced Driver Assistance Systems. Therefore, it is crucial to address these challenges to ensure the effective operation of systems under low-light conditions.
The problems of low-light image enhancement (LLIE) are commonly addressed using two main methods: Histogram Equalization [1] and Retinex [2]. Histogram Equalization works by enhancing contrast through the redistribution of grayscale values. On the other hand, the Retinex theory divides the image into reflectance and illumination components to improve reflectance and overall image quality. Techniques such as Single Scale Retinex [3] and Multi-Scale Retinex [4] are particularly effective in preserving details and managing complex lighting conditions.
With technological advancements, various deep learning-based methods [5, 6, 7] have been proposed to enhance the quality of low-light images. However, these methods often require substantial computational resources, which limits their practical application on real-world devices. Therefore, designing lightweight and efficient image enhancement techniques is crucial. In our previous study, we introduced the CPGA-Net [8], which combines Retinex theory with the Atmospheric Scattering Model (ATSM) [9] and utilizes gamma correction for both global and local processing; it highlights the importance of gamma correction in LLIE. CPGA-DIA [10] explores exposure correction and low-light enhancement issues through dynamic gamma adjustment, showing that gamma correction can be efficient and effective for the enhancement process even in deep learning frameworks.
Figure 1: Performance Comparison with SOTA approaches. (a) Comparison of PSNR vs. Number of Parameters on LOLv1; (b) Comparison of mean NIQE vs. Number of Parameters on unpaired data.
Building on these ideas with a theoretical-based structure and gamma-correction prior, we expanded CPGA-Net by rethinking the theoretical equation as an attention mechanism and transitioning the gamma estimation module from global processing to local processing. Our main contributions are as follows:
  • Extended CPGA-Net: We propose an enhanced version of the Channel-Prior and Gamma-Estimation Neural Network (CPGA-Net) called CPGA-Net+. This model achieves state-of-the-art (SOTA) image quality and efficiency performance for both supervised and unsupervised learning, making it a lightweight and practical solution for real-world applications, as shown in Fig. 1.
  • Theoretical-based Attention for Illumination: We modularized the Atmospheric Scattering Model into a block design and incorporated gamma correction into the local branch. This architecture modification significantly improves image structure and detail, maximizing the efficiency of prior knowledge for illumination to strengthen the overall image quality.
  • See Through the Mechanism Behind the Image Enhancement: We evaluate our approach through the degradation formula and optimization, demonstrating that the optimal solutions align with the output features from the neural network, ensuring strong interpretability and robustness.
  • Block Design Simplification for Lightweight Implementation: We can directly remove the local branch of our network during inference, which induces a slight performance drop but significantly improves efficiency, reducing the number of parameters and FLOPS by about 66.8% and 77.1%, respectively.
II Related Work
Our work enhances the original purely convolutional architecture by integrating an attention mechanism while preserving the benefits of a lightweight and efficient design. This advancement is particularly well-suited for LLIE tasks. This section will conduct a comprehensive literature review on leveraging a deep learning-based approach in LLIE and explore developments in lightweight model architectures.
II-A Deep Learning-Based LLIE
With the continuous development of LLIE, Retinex theory has increasingly demonstrated its potential in conjunction with deep learning techniques. Several methods based on this approach address low-light environments. For instance, while LIME [11] differs from directly decomposing images according to Retinex theory, it primarily relies on estimating the illumination map of low-light images for enhancement. RetinexNet [5] and KinD [7] decompose images into reflectance and illumination components during decomposition. In the adjustment phase, they adjust the illumination component’s brightness and denoise the reflectance component, ultimately merging them based on the theory to restore natural images. EnlightenGAN [6] proposes an unsupervised Generative Adversarial Network (GAN) that can be trained without paired low/normal light images. It introduces a global-local discriminator structure, self-regularized perceptual loss fusion, and attention mechanisms to enhance image quality. LLFlow [12] presents a flow-based approach for LLIE by modeling distributions of normally exposed images. It improves traditional methods by using an illumination-invariant color map as the prior distribution rather than a Gaussian distribution. The process features an encoder to extract stable color attributes and an invertible network to map low-light images to normally exposed image distributions, aiming for improved enhancement performance.
II-B Lightweight LLIE
In the context of LLIE, developing lightweight methods is crucial for practical deployment, often requiring sophisticated techniques to achieve both efficiency and effectiveness. For example, Zero-DCE [13] replaces the direct image enhancement process with a curve-fitting approach, introducing a series of reference-free loss functions that reduce the computational burden, achieving an efficient and lightweight design. RUAS [14] builds upon Retinex theory by proposing a Retinex-inspired model that leverages prior information from low-light images, combined with a distillation unit-based search architecture and a cooperative bilevel search strategy, maintaining high performance while achieving a lightweight design. IAT [15] decomposes the task into local and global processing components. The local branch leverages a convolution-based Transformer to perform image restoration and enhancement. In contrast, the global branch utilizes global priors, including color transformation matrices and gamma correction, to apply global adjustments across different exposure conditions, thereby attaining efficient and lightweight performance improvements. PairLIE [16] deviates from the traditional Retinex approach of directly decomposing images; instead, it removes noise through a self-supervised mechanism before decomposition. It shows that training on low-light images of the same scene with different exposures better learns features. Finally, it merges them using a simple convolutional network to achieve a lightweight design. Inspired by Retinex theory and ISP (Image Signal Processor) frameworks, FLIGHT-Net [17] features Scene Dependent Illumination Adjustment for illumination and gain processing and Global ISP Network Block for compact color correction and denoising. This design optimizes for both efficiency and lightweight operation.
II-C Insights and Innovations
Our proposed method builds on two key insights from prior work: theory-driven attention mechanisms and lightweight design principles in LLIE. Our approach leverages the theoretical basis of perceiving visual information through air turbulence—a principle effectively utilized in Retinex-based [5, 6, 7, 11, 12, 14, 16] and ATSM methods [8, 10, 18]. We incorporate this concept into a streamlined attention module that selectively enhances key features, guided by prior knowledge [8, 13, 19], to improve detail preservation and contrast. Furthermore, following [8, 10, 15], gamma correction is integrated into our model and has extended its application from global to local processing to maximize adaptability via environmental characteristics. These attention modules and formulae turn our model apart by integrating theoretical principles with practical design, delivering high-quality enhancement while preserving a lightweight architecture.
III Methodology
In this section, we delve into the reconstruction of ideal images by global and local concepts in image processing, leveraging advanced deep learning techniques. The discussion will commence with the theoretical underpinnings and motivations for developing CPGA-Net+, followed by an exposition of the network’s architecture and implementation, which can be separated as Atmospheric Scattering-driven Attention and Plug-in Attention with Gamma Correction, as shown in Fig. 2.
Figure 2: The schematic of our proposed approach, CPGA-Net+.
III-A The Connection Between Theoretical Equations and Low-light Image Enhancement
Guo et al. [11] figured out that the Retinex theory [2] and the ATSM [9] have a strong relationship that can represent each other by reformulating their equations. Retinex theory assumes that the received image can be decomposed into illumination and reflectance components, as shown in Eq. (1):
$S=i\cdot R$ (1)
where $S$ is the perceived image, $i$ denotes the illumination component, and $R$ represents the reflectance component.
On the other hand, the Atmospheric Scattering Model for haze removal is defined as:
$I=tJ+(1-t)A$ (2)
where $I$ represents the input image, $J$ represents the haze-free image, $t$ represents the atmospheric transmission, and $A$ represents the intensity of atmospheric light.
Based on Dong et al. [18], the low light image $I$ can be seen as $1-L$, where $L$ represents the low-light image, and $J$ can be seen as $1-R$, where $R$ reflects the important characteristics of the input image. The above substitutions are performed under the normalized pixel values in [0, 1]. With these substitutions, we rewrite Eq. (2) into the following form:
$R=\tilde{t}L+(1-\tilde{t})\tilde{A}$ (3)
where $\tilde{A}=1-A$ and $\tilde{t}=1/t$. The model Eq. (3) described is the cornerstone of our neural network design. Part of the information of the reflectance $R$ comes from the known image $L$, part of it comes from an unknown image $\tilde{A}$, and the proportion sum of their contributions is limited to 1. When $L$ is very dark or noisy (the scene information is less reliable), the contribution of $L$ is lowered, and the contribution of $\tilde{A}$ is increased; when $L$ is relatively bright and less noisy (the scene information is more reliable), the contribution of $L$ is increased, and the contribution of $\tilde{A}$ is reduced. So, $\tilde{t}$ should reflect the intensity level of $L$ in some way.
This reformulation reveals an alternative imaging perspective, where $L$ is linked to the characteristic of atmospheric light, which predominantly includes environmental interference, $\tilde{A}$ corresponds to the unknown noise-free image, and $R$ corresponds to the reflectance in a linear relationship between $L$ and $\tilde{A}$. The underlying mechanism of this formulation matches the phenomenon of our atmospheric scattering-based approach and will be discussed in Section III-D.
In our previous work, CPGA-Net [8], we successfully utilized these theoretical equations in deep learning form. In this article, we extended the idea and proposed an attention mechanism called the Channel-Prior block (CP block), which modularizes the relationship of both equations into a systematic form. We restructured the module with convolutions and a ResBlock while fusing features with the original RGB channels at each step to streamline the module into an attention-block design. This helps in activating the feature map to align with the original channels. In CPGA-Net, three-channel priors are selected as the input for $\tilde{t}$ estimation: the Bright Channel Prior (BCP), the Dark Channel Prior (DCP), and the luminance channel (Y component from the YCbCr color space). They can be defined respectively as:
$I^{\text{bright}}=\max_{c\in\{r,g,b\}}\left(I^{c}\right)$ (4)
$I^{\text{dark}}=\min_{c\in\{r,g,b\}}\left(I^{c}\right)$ (5)
$I^{\text{luminance}}=0.299\cdot I^{r}+0.584\cdot I^{g}+0.114\cdot I^{b}$ (6)
where $I^{c}$ represents the color channel $c$ of the input image $I$. These common features represent the brightness variation in different environments and are widely used in traditional methods [19]. The combination of channel priors shows sensitivity to contrast, which is an important clue in representing the atmospheric transmittance $\tilde{t}$ to guide the enhancement, as shown in Fig. 3.
Figure 3: The comparison between the brightness representation. From top to bottom are the original image, edge detection of the luminance channel $I^{\text{luminance}}$, and edge detection of the channel-prior which is $\max(I^{\text{bright}},I^{\text{dark}},I^{\text{luminance}})$.
Additionally, to transform the priors into a high-dimensional processing module rather than maintaining the original channels, we simplified the luminance channel into a more basic representation—specifically, the mean of the channels, akin to the concept of intensity (such as the I component in the HSI color space), which also indicates image brightness. Therefore, the channel priors are depicted as:
[…]
where $f$ is the input feature map, and $F^{\text{CP}}$ denotes the channel prior features. If $c$ consists of RGB channels, $f$ will equal the $I$ that appears in Eq. (4), (5), and (6) As a result, the channel priors are simplified to the input channel’s max, min, and mean, making the attention module more responsive to the overall contrast control of the image.
For the $\tilde{A}$ estimation, which captures detailed features and reconstructs the image, we redesigned it as a mini-U-Net-based architecture with encoder and decoder pathways connected by skip connections. This design effectively captures and preserves spatial information at multiple scales, enhancing the model’s ability to reconstruct fine details in the image—a technique commonly used in LLIE, such as [5, 6, 7]. Due to considerations for lightweight efficiency, we only downsample the input once, reducing computational complexity while maintaining adequate feature extraction capabilities. This approach ensures that the model remains efficient and suitable for real-time applications or scenarios with limited computational resources without significantly compromising the quality of the reconstructed image.
After obtaining the estimates of $\tilde{t}$ and $\tilde{A}$, we can reconstruct our features using Eq. (3), which serves as an attention module sensitive to brightness variations in the scene. This leads to the proposed Atmospheric Scattering-driven Attention, formulated as follows:
$R_{\text{att}}(f)=\tilde{t}(f,F^{\text{CP}})L^{\prime}(f)+\big{[}1-\tilde{t}(f,F^{\text{CP}})\big{]}\tilde{A}(f)$ (8)
where $L^{{}^{\prime}}$ is the mapped input tensor with a matching channel for formula calculation, $\tilde{t}(f,F^{\text{CP}})$ indicates the derived transmittance with input feature map $f$ and channel-prior features $F^{\text{CP}}$, and the structure is shown in Fig. 4.
Figure 4: The block diagram of the Channel-Prior Block (CP block), where c denotes a concatenation operation.
III-B Relinking the Gamma Correction to Global-Local Processing
Not only taking advantage of the characteristics of low-light images, CPGA-Net [8] also combines gamma correction based on IAT [15]. Gamma correction is a simple technique that adjusts all the pixels with pointwise exponential operations, as shown in Eq. (9):
$s=r^{\gamma}$ (9)
where $\gamma$ is the gamma value controlling the degree of correction, enhancing the input image $r$ to produce the output image $s$. These approaches combined an independent branch of regression with the enhancement model to better estimate gamma values. Moreover, the complexity of gamma value estimation can complicate training goals and make the process prone to divergence. Taking these aspects into account, we proposed an IAAF (Intersection-Aware Adaptive Fusion) module, as shown in Fig. 5 and Eq. (10):
$\begin{split}\hat{R}&=\text{IAAF}(R^{\gamma},R)=(R\cup R^{\gamma})-(R\cap R^{\gamma})\\ &\approx R+R^{\gamma}-\cap(R,R^{\gamma})\end{split}$ (10)
where the enhanced image $\hat{R}$ is created by combining $R$ and $R^{\gamma}$ while removing any overlapping elements and $\cap(R,R^{\gamma})$ represents the intersection estimation for finding similar features across gamma-corrected and uncorrected images.
Figure 5: A block diagram of the IAAF (Intersection-Aware Adaptive Fusion) module [8].
First, based on the findings from [20, 21], we know that downsampling the image resolution by a limited scale does not significantly impact performance while greatly reducing computational cost. Therefore, we redesigned the process of gamma value estimation, as shown in Fig. 2. We reduced the feature resolution in the first ResCBAM block by adjusting the stride from 1 to 2. This adjustment enables subsequent computations to process half-sized images while leveraging residual and downsampling connections. This design improves the learnability of the IAAF module, facilitating more effective integration of global and local processing.
Furthermore, we adapt this operation to function within the Channel-Prior Block by incorporating it directly into the attention process, allowing it to act as a global feature that estimates the optimal gamma-correction value for each feature channel rather than applying gamma correction only at the final step as commonly done in [8, 15]. The modified attention output can then be expressed as:
$\hat{R}_{\text{att}}=\text{IAAF}(R_{\text{att}}^{\gamma},R_{\text{att}})+R_{\text{att}}$ (11)
where $\hat{R}_{\text{att}}$ is the attention feature built on the gamma corrected feature $R_{\text{att}}^{\gamma}$ and the uncorrected feature $R_{\text{att}}$. Moreover, we add another residual to this attention in Eq. (11), distinguishing it from the reconstruction applied in Eq. (10), making it an auxiliary attention to support the residual features.
Building on insights from [10], we understand that the adaptive gamma value serves as an environmental factor representing the overall illumination conditions, which can vary across different scenes. Consequently, this attention module ensures that the network focuses on regions within the broader image where gamma correction yields the most significant enhancement. This approach is called “Plug-in Attention with Gamma Correction,” and the CP block with IAAF is named “CPGA block.”
III-C Loss Functions
In supervised learning, we use four loss functions to guide our approach: L1 loss, perceptual loss, HDR L1 loss, and SSIM loss. We utilize HDR L1 loss, perceptual loss, and total variation loss for unsupervised learning, with the target set as the histogram-equalized low-light image.
The L1 loss function, a commonly use loss function that performs better in image enhancement and restoration, is defined as:
$L_{1}=\|\hat{Y}-Y^{\text{GT}}\|_{1}$ (12)
where $\hat{Y}$ is the output and $Y^{\text{GT}}$ is the ground truth.
Perceptual loss [22] is commonly used in image restoration, style transfer, and generation. It emphasizes capturing high-level features and structures that closely resemble human perception. The loss can be expressed as:
$L_{per}=\|\Psi(\hat{Y})-\Psi(Y^{\text{GT}})\|_{2}^{2}$ (13)
where $\Psi$ represents the feature extractor of VGG16.
HDR L1 loss, as introduced in [23] is computed in the tone-mapped domain since HDR (High Dynamic Range) images are typically viewed after tone-mapping. To achieve this, they apply the widely used µ-law function to calculate the loss:
$T=\text{sgn}\frac{\log(1+\mu x)}{\log(1+\mu)}$ (14)
where $\mu$ is set to 5000, $T$ is the tone-mapped HDR image, and $x$ is the input image. Then, the $\mu$-law function is applied to the L1 loss as follows:
$L_{\text{HDR-L1}}=\|T(\hat{Y})-T(Y^{\text{GT}})\|_{1}$ (15)
It represents the image in the tone-mapped domain, ensuring the loss is calculated in a perceptually relevant space that aligns with how normal light images are typically viewed.
SSIM loss is a function that measures the similarity between two images based on structural information via the SSIM index (structural similarity index). It compares luminance, contrast, and structure, reflecting perceptual quality better than traditional pixel-wise losses. It can be written as:
$L_{\text{SSIM}}=1-\text{SSIM}(\hat{Y},Y^{\text{GT}})$ (16)
For unsupervised learning of our approach, the only optimal for brightening is the histogram-equalized low-light image, which is a good aim for the contrast but not for the details. Thus, we utilized total variation loss [24, 25] for denoising and improving the smoothness:
$L_{TV}=\frac{1}{hwc}\sum_{i,j,k}\sqrt{\left(\hat{Y}_{i,j+1,k}-\hat{Y}_{i,j,k}\right)^{2}+\left(\hat{Y}_{i+1,j,k}-\hat{Y}_{i,j,k}\right)^{2}}$ (17)
where $h$, $w$, and $c$ represent the height, width, and number of channels, respectively; $i$, $j$, and $k$ represent the indices corresponding to height, width, and channel, respectively.
III-D Revealing the Explainable Mechanisms Behind Degradation
Our method follows a rule-based learning strategy, leveraging the ATSM and the Retinex theory to perform image processing based on physical models. The ATSM simulates the scattering and absorption of light in uneven media, while the Retinex theory, inspired by retinal imaging, models the human visual system. Although these methods provide solid theoretical support from both physical and physiological perspectives, challenges remain in the interpretability of neural networks. While applying the models to deep learning, it is difficult to explain its decision-making mechanisms due to the highly nonlinear nature of deep learning models. Similarly, these are hard to define and are barely explained through experience or a simple understanding of the equations from the generated features in the neural network. To improve the interpretability of our approach, we utilize the traditional image degradation model, which offers a more explicit theoretical framework for the deep learning process. This effectively enhances the transparency of the entire neural network, opening the “black box” of image enhancement.
Initially, the fundamental image degradation model in matrix form [1] can be defined as:
[…]
where G is the image after degradation, H is the unknown degradation kernel, N denotes additive noise, and F is the image before degradation. The restoration process is aimed at estimating F by:
$\hat{\textbf{F}}=\textbf{H}^{-1}(\textbf{G}-\textbf{N})=\textbf{H}^{-1}\textbf{R}$ (19)
[…]
Figure 6: Visualization of extracted features corresponding to Eq. (3) and (11) [8]. (a) Low-light image $L$; (b) a portion of $L$; (c) $\tilde{t}$ (contribution proportion from $L$); (d) An estimated image of $\tilde{A}$; (e) Reconstructed low-light image $R$; (f) Gamma-corrected image $R^{\gamma}$; (g) Enhanced image $\hat{R}$.
We attempt to further verify our hypotheses for the local branch through least-squares optimization [26]. First, we make the following simple assumptions: we assume that $\hat{A}$ and $L$ can be expressed as the noise-free reconstructed image $R^{\text{GT}}$ with additional composite noises that consist of three-channel mixing factors, lighting changes, and inherent thermal noises, which are particularly prominent in low-light environments, as depicted below:
$\tilde{A}=R^{\text{GT}}+N_{A}$ (20)
$L=R^{\text{GT}}+N_{L}$ (21)
where $L$ is the low-light image, $N_{A}$ and $N_{L}$ denote the noise components associated with $\tilde{A}$ and $L$, respectively. While estimating $N_{A}$, we assume $N_{L}$ is fixed once the low-light image $L$ is given. Then, we can define the following residual error $r(\tilde{t},N_{A})$ by combining Eq. (3), Eq. (20), and Eq. (21):
$r(\tilde{t},N_{A})=R-R^{\text{GT}}=(1-\tilde{t})N_{A}+\tilde{t}N_{L}$ (22)
where $R$ is the reconstructed image, $\tilde{t}$ is a positive parameter used to adjust the reconstruction process. We aim to find $N_{A}$ and $\tilde{t}$ such that the residual error is minimized, allowing $R$ to approach $R^{\text{GT}}$.
We further define a cost function $C(\tilde{t},N_{A})$ for the least-square optimization process:
$C(\tilde{t},N_{A})=\frac{1}{2}\left[r(\tilde{t},N_{A})\right]^{2}$ (23)
Next, to find the stationary points in the optimization process, we conduct partial differentiation of $C(\tilde{t},N_{A})$ with respect to $\tilde{t}$ and $N_{A}$:
$\frac{\partial C}{\partial\tilde{t}}=r(\tilde{t},N_{A})\cdot\left({N_{L}-N_{A}}\right)=0$ (24)
$\frac{\partial C}{\partial N_{A}}=r(\tilde{t},N_{A})\cdot\left(-\tilde{t}\right)=0$ (25)
Case I: $r(\tilde{t},N_{A})\neq 0$: We have
$(N_{A}-N_{L})=0\quad\text{and}\quad\tilde{t}=0$ (26)
which leads to the solution $\tilde{t}=1$ and $N_{L}=N_{A}$, resulting in:
$r_{\text{min}}(\tilde{t},N_{A})=N_{A}$ (27)
However, $\tilde{t}=0$ results in $R=L$ from Eq. (3), meaning that no restoration effect is involved at all (identity mapping from the input $L$ to the output $R$).
Case II: $r(\tilde{t},N_{A})=0$: We have
$R-R^{\text{GT}}=0\implies R=R^{\text{GT}}$ (28)
and
$(1-\tilde{t})N_{A}+\tilde{t}N_{L}=0\implies N_{L}=\left(1-\frac{1}{\tilde{t}}\right)N_{A},\quad\tilde{t}\neq 0$ (29)
Therefore, the optimal solution exists when Eq. (29) holds.
The results show that the enhancement follows the degradation formula, with the ATSM simulating the network’s processing of illumination, reflection, and noise, supporting the rationality of neural networks in LLIE. This enhances the model’s transparency and reveals its interpretability mechanism, offering deeper insights into how neural networks function in image enhancement tasks.
From a learning mechanism perspective, deep learning exhibits significant similarities to human cognition, emphasizing that the most prominent features carry essential information rather than relying on intricate details, with $\tilde{A}$ for detail restoration representing more critical information than $\tilde{t}$ for contrast and saturation. By modularizing and extending the understanding through attention mechanisms, we effectively capture the relationships between local and global information, enhancing image quality in low-light conditions and yielding better results across various complex scenarios. This method underscores the potential for integrating traditional theory with deep learning models, offering valuable insights for future technological advancements and opening new research opportunities in image enhancement.
IV Experiment Results
This section compares our approach with several SOTA methods on benchmark datasets, including paired and unpaired datasets.
IV-A Datasets and Evaluation Metrics
For evaluation, we apply our approach to both paired and unpaired datasets. For paired data, we use the LOLv1 and LOLv2 datasets [5], benchmarks for the LLIE task. LOLv1 includes 485 images for training and testing, while LOLv2 consists of two subsets: real-captured and synthetic. The real-captured subset (LOLv2 Real) has 689 images for training and 100 for testing, while the synthetic subset (LOLv2 Synthetic) has 900 training images and 100 testing images. For unpaired data, we utilize five datasets: LIME [11], MEF [27], NPE [28], VV [29], and DICM [30]. Since these datasets lack ground truth references for paired evaluation, we assess performance using the NIQE metric, which is widely used to evaluate the naturalness of images. For object detection applications, we apply our approach to the Exclusively Dark (ExDark) dataset [31], which consists of 7,363 low-light images from very low-light environments to twilight with 12 object classes and provides a suitable benchmark for evaluating object detection performance in such conditions.
IV-B Implementation Details
Our experiments are conducted and evaluated with an NVIDIA GeForce RTX 3090 GPU. For the preprocessing of training, we cropped the image into $256\times 256$ pixels with random translation. We set 600 epochs for training. An initial learning rate of $10^{-3}$is set for training with the Adam optimization scheme, and the learning rate changing cycle is 67 epochs by the Cosine Annealing scheduler. To address computational demand, we implemented simplified block pruning named “Block Design Simplification (BDSF)” in our lightweight version, directly removing blocks in the local branch for inference. Furthermore, we have also applied our approach as unsupervised learning using histogram equalization, demonstrating its flexibility. When using unsupervised learning, we trained on LOLv2 Real for 50 epochs, which consists of more realistic data and does not require any normal exposure images for training.
IV-C Evaluation Results
As shown in Tables I and II, our approach reaches a high standard compared to others, which ranked third on paired datasets and ranked first on five sets of unpaired data, while the model remains lightweight design with low numbers of parameters and FLOPs. The visualized comparisons are shown in Figs. 7 and 8. Also, we figured out that the approaches with CPGA architecture present better quality on unpaired data, which means that our theoretical equations assumptions for improvement are closer and related to nature, making the image more realistic. Furthermore, our method successfully improves the performance via the same architecture by 5% SSIM compared to the CPGA-Net while maintaining the lightweight design.
For the comparison of unsupervised approaches, as shown in Table II we ranked first compared to other unsupervised approaches with better contrast. This demonstrates the robustness of our theoretical-based network architecture when using simple supervision of histogram-equalized images. However, there are more noticeable defects and distortions due to the lack of strong supervision of the details, as shown in Fig. 9. This will be a focus for our future work.
TABLE I: Comparison to SOTA methods on paired datasets [5]. We represent the first and second ranks with bold and underlined, respectively. BDSF means Block Design Simplification for our approach.
LOLv1 LOLv2-real LOLv2-syn Efficiency
[…]
LIME [11] 16.67 0.560 0.368 15.24 0.470 17.63 0.787 - -
Retinex-Net [5] 16.77 0.425 0.474 18.37 0.723 17.14 0.756 0.555 587.470
KinD [7] 17.65 0.771 0.175 14.74 0.641 17.28 0.758 8.160 574.950
EnGAN [6] 17.54 0.664 0.326 18.23 0.617 16.49 0.771 114.350 223.430
Zero-DCE [13] 14.86 0.562 0.335 14.32 0.511 17.76 0.814 0.075 4.830
RUAS [14] 18.23 0.720 0.270 15.33 0.488 13.76 0.634 0.003 0.830
IAT [15] 23.38 0.809 0.210 23.50 0.824 15.37 0.710 0.091 5.271
PairLIE [16] 19.56 0.730 0.248 19.89 0.778 19.07 0.794 0.342 81.838
FLIGHT-Net [17] 24.96 0.850 0.134 21.71 0.834 24.92 0.930 0.025 3.395
DDNet [32] 21.82 0.798 0.186 23.02 0.834 24.63 0.917 5.390 111.47
CPGA-Net [8] 20.94 0.748 0.260 20.79 0.759 20.68 0.833 0.025 6.030
CPGA-DIA [10] 20.37 0.760 0.280 22.18 0.794 18.22 0.799 0.065 15.520
CPGA-Net+ 22.53 0.812 0.205 20.90 0.800 23.07 0.907 0.060 9.356
CPGA-Net+ (BDSF) 22.53 0.812 0.205 20.90 0.800 23.07 0.907 0.020 2.141
TABLE II: The image quality comparison on unpaired data [11, 27, 28, 29, 30] in terms of the NIQE metric, where lower values generally indicate better performance. We represent the first and second ranks with bold and underlined, respectively. For the learning methods, T indicates the traditional approach, U indicates unsupervised learning, and S indicates supervised learning. BDSF means Block Design Simplification for our approach.
Datasets
Original Image and Method Types MEF LIME NPE VV DICM Avg
Low-light Image N/A 4.2650 4.4380 4.3190 3.5350 4.2550 4.1624
NPE [28] T 3.5240 3.9048 3.9530 2.5240 3.7600 3.5332
LIME [11] T 3.7200 4.1550 4.2680 2.4890 3.8460 3.6956
EnlightenGAN [7] U 3.2320 3.7190 4.1130 2.5810 3.5700 3.4430
Zero-DCE [13] U 4.0410 3.7890 3.5041 2.7526 3.1018 3.4377
RUAS [14] U 4.1403 4.2900 4.8713 3.5086 4.5417 4.2704
PairLIE [16] U 4.0862 4.3113 4.0890 3.1595 3.2422 3.7776
CPGA-Net+ [8] U 3.5950 3.2575 3.4438 2.8820 3.0350 3.2427
KinD [7] S 3.8830 3.3430 3.7240 2.3208 2.9888 3.2519
IAT [15] S 3.6188 4.1722 3.2890 2.5270 3.0325 3.3279
FLIGHT-Net [17] S 3.5491 3.7049 3.3311 2.9435 2.8979 3.2853
DDNet [32] S 3.2734 3.4329 3.1135 2.0223 2.6409 2.8970
CPGA-Net [8] S 3.8698 3.7068 3.5476 2.2641 2.6934 3.2163
CPGA-DIA [10] S 3.5880 3.5570 3.1650 2.0930 2.6300 3.0006
CPGA-Net+ S 3.4968 3.0626 3.0886 1.9133 2.8282 2.8779
CPGA-Net+ (BDSF) S 3.4969 3.0655 3.0881 1.9136 2.8268 2.8782
Figure 7: Visual comparison on paired datasets [5].
Figure 8: Visual comparison on unpaired datasets [11, 27, 30].
Figure 9: Visual comparison of unsupervised approaches on unpaired data [28, 29].
IV-D High-Level Vision Task
In this section, we address the challenge of objection detection in low-light environments by utilizing a joint training approach of YOLOv9s [33] with the SOTA approaches of LLIE on the ExDark dataset [31], as illustrated in Table III. Our approach improves the mean Average Precision (mAP) by 0.075 compared to baseline. All the LLIE methods listed here can improve object detection performance, among which the Zero-DCE [13] and our proposed method are the best. The results show that our proposed method can improve not just human perception but machine perception as well.
TABLE III: Comparison of performance metrics between YOLOv9s with CPGA-Net+ and other SOTA methods on the ExDark dataset [31].
[…]
V Ablation Study
In this section, we analyze the effectiveness of each systematic module and training technique, including the systematic design, the number of Channel-Prior blocks, and loss functions.
V-A Systematic Design and Integration
As shown in Table IV, our method effectively fuses the gamma correction from the global branch to the local branch, resulting in improved overall performance and demonstrating the strength of our approach. By grounding the attention mechanism in gamma correction, we ensure that the enhancement process remains aligned with the non-linearities inherent in both the imaging process and human perception.
On the other hand, we also figured out that the global branch alone performs surprisingly well. In Section III-D, we elaborate on the system underlying the equations and highlight the relationship between $R$ and $R^{\text{GT}}$, and the results suggest that this correlation is stronger than expected. For instance, Table IV(c) and Table IV(e) show only a minor difference in PSNR (0.5 dB), and there is no performance difference between Table IV(e) and Table IV(f) on LOLv1. However, despite the significant contribution of the global branch, Table IV(c) still shows a noticeable gap in achieving optimal performance without incorporating local processing. This underscores the necessity of the local branch to bridge the gap between local and global processing, ensuring the system achieves its highest potential in quality and consistency. Since our pruning approach is based on the concept of block pruning but differs from existing methods [34, 35, 36, 37] and retains its simplicity even without requiring additional training, we named the proposed technique ”Block Design Simplification (BDSF).”
TABLE IV: Ablation study of systematic design. L-G denotes our design of plug-in attention from global to local processing, utilizing the CPGA block to bridge the gap between local and global branches. (f) shares the weights from (e) but performs inference via global processing only, sharing the same design as (c). (e)* denotes using the weights obtained from (e).
[…]
V-B The Numbers of Channel-Prior Blocks
We explore how varying the number of CP blocks affects the model’s capacity to enhance image quality. The results, summarized in Table V, show that increasing the number of CP blocks leads to an improvement from 0 to 2 blocks but show no significant changes from 2 to 4 blocks. However, both the number of parameters and computational cost (FLOPs) increase with more CP blocks, introducing greater computational demands. Therefore, the optimal number of CP blocks should balance performance gains with resource efficiency. For our final approach, we selected 2 CP blocks to achieve a lightweight and efficient design.
TABLE V: Ablation study of the number of CP blocks. $N=2$ is the default setting of our approach.
[…]
0 20.56 0.754 0.250 0.034 4.243
2 22.53 0.812 0.205 0.060 9.356
4 21.93 0.805 0.215 0.087 14.503
V-C Loss Functions
This section examines the impact of various loss function combinations on the model’s performance. We tested L1 loss, Perceptual loss, HDR L1 loss, and SSIM loss, with the results summarized in Table VI. Using the default settings as in CPGA-Net, the combination of L1 and Perceptual losses performs well, yielding a PSNR improvement of 2.82 dB and an SSIM increase of 0.024. The HDR L1 loss significantly enhances all three metrics, with PSNR increasing by 2.86 dB, SSIM by 0.041, and LPIPS decreasing by 0.085. While SSIM loss improves its specific metric with an SSIM boost of 0.031, it is less effective in enhancing PSNR. Ultimately, combining all these losses results in the best overall performance for supervision, which improves PSNR by 4.04 dB, SSIM by 0.083, and LPIPS by 0.118.
TABLE VI: Ablation study of loss functions.
[…]
VI Limitation
While our proposed method shows notable improvements in efficiency and performance, a key limitation warrants consideration. As discussed in the methodology section, our approach benefits from the guidance of Channel-Priors and Gamma Correction, which enhances contrast and visual perception. However, gamma correction for brightening with a small value of estimated gamma can lead to defects or distortions in extreme low-light scenarios, even while maintaining realistic exposure. Fig. 10 shows an example of such a limitation. Addressing this issue will require further refinement for brightening and denoising, which we aim to pursue in future work.
Figure 10: Visualization of extreme low-light scenarios [11, 30]. (a) Original images; (b) Supervised CPGA-Net+; (c) Unsupervised CPGA-Net+. While the enhancement tone appears satisfactory, both methods exhibit defects such as pepper-and-salt noise, resulting in a grainy texture.
VII Conclusion
This work looks deeper at CPGA-Net, utilizing it as an attention mechanism grounded in theoretical formulas. We propose a stacked and modularized attention module to focus on image details. Additionally, we integrate gamma correction into the local branch, creating a Plug-in Attention module for each CP block. This enhancement makes our approach lightweight yet SOTA in performance, maintaining strong efficiency and stable operation across devices with limited computational resources. In the future, we are striving to improve the unsupervised learning of CPGA-Net+ and integrate our approach into HDR imaging and exposure fusion, improving detail preservation in both bright and dark areas by leveraging brightness sensitivity through prior knowledge, such as channel and gamma-correction priors. This will enhance the output’s dynamic range and overall fidelity in real-world applications.
Acknowledgment
This work was supported in part by the Ministry of Science and Technology, Taiwan, under Grant NSTC 113-2221-E-033-055.